# Learning interactions between Rydberg atoms

### v1

@@ III.2.2 Effect of $\vec{\Omega}$-histories
III.2.3 Number of physical system sizes in training datasets
@@
To determine the effect of the number of physical system sizes included in the training dataset, we plot in Fig. 7, for the set of sizes $\{4\times 4,5\times 5\}$ and $\{4\times 4,5\times 5,6\times 6\}$, the coefficient of determination $R^{2}$ (16) in the top panel, the MAE (17) in the middle panel, and the median of the MAE (MEDAE) in the bottom panel. The case #$\#3$ is utilized. Without much surprise, the GNN performs much better at extrapolating relative NN distances $\Delta R_{\langle i,j\rangle}$ for clusters whose sizes lie outside the training set. It may, however, come as a surprise that the training set $\{4\times 4,5\times 5\}$ performs better than that having $\{4\times 4,5\times 5,6\times 6\}$ for the common training sizes, but it seems that this effect comes from the inclusion of the $\chi^{\text{NNN}}$, as seen in Fig. 4. Since smaller clusters are subparts of larger clusters, the information of the smaller clusters aggregates and makes it possible to predict with high precision the $\Delta R$’s of clusters with sizes beyond the training set.
Figure 8: Target optimization schemes. Incidence on various metrics of including the NNN relative distance as a target in the cost function minimization. In blue and pink, both the NN and NNN spin correlation functions are employed to build the edge feature space, while in red and violet, only the NN ones are used. Training sizes used are indicated in parentheses. The vertical grey lines refer to the point beyond which sizes lie outside of the training range.
III.2.4 Predicting NN and NNN distances
@@
Now, regarding the cost function in Eq. (13), one could ask whether one should use both the relative NN and NNN distances whenever both $\chi^{\text{NN}}$ and $\chi^{\text{NNN}}$ are provided, or if one should use an overcomplete set of physical short-range correlators to predict only the relative NN distances, i.e. $\chi^{\text{NN}}$ and $\chi^{\text{NNN}}$. As proven in Sec. II.2, with arbitrary precision on the correlation functions, one would theoretically only need $\chi^{\text{NN}}$ to predict relative NN distances $\Delta R_{\langle i,j\rangle}$ between Rydberg atoms. However, we may think that giving some extra information to the GNN could be beneficial to overcome numerical or experimental imprecision. To that effect, in Fig. 8, we compare two GNN training cases #$\#3$ where, on one hand, only relative NN distances are used as targets, and on the other hand, both NN and NNN relative distances are used as targets, for training sizes $\{4\times 4,5\times 5,6\times 6\}$ and $\{4\times 4,5\times 5,6\times 6,7\times 7\}$. It is very clear that the overcompletion of information helps the GNN palliate the numerical/experimental inaccuracies in determining the relative distances, meaning that predicting solely NN relative distances with $\chi^{\text{NN}}$ and $\chi^{\text{NNN}}$ at hand is better than predicting both NN and NNN relative distances with the same information throughput. It also seems that predictions of larger relative distances (NNN and above) require larger cluster sizes in the training set because edge effects prevail longer. Therefore, if one seeks to also predict NNN relative distances between atoms, one should consider training the GNN with larger cluster sizes to improve substantially the predictive power of the model (compare the navy blue curve with the pink one in Fig. 8). Interestingly, adding another task network specializing in predicting NNN distances worsened the GNN performance compared to using the same task network for predicting NN and NNN distances.
III.3 Predictions from snapshots
In this section, we ask how the GNN performs when the input correlation functions are estimated from a collection of successive projective measurements of the wave function instead of computed exactly. Note that all the assertions made in the previous sections carry over to this training since the GNN does not change, nor does the physics. Here, we still train the GNN with a training dataset whose samples are made up of correlation functions (local and nonlocal) estimated from a number of snapshots (projective measurements of the wave function), nicknamed ‘snapshot data’. One could naively think that predicting the NN relative distances of ‘snapshot data’ with the GNNs trained with ‘exact’ correlators (directly measured in the ground state) would lead to a better outcome as opposed to the case where the GNNs are trained with snapshot data. We tried predicting the relative distances from snapshot data using a GNN model trained with ‘exact’ correlators, but it produced worse results compared to a GNN model trained on the snapshot data (not shown). In fact, when measuring directly from the wave function, no errors coming from the finite number of projective measurements spoil the training dataset, and the GNN can thereof dedicate all its variational parameters to ‘learn’ the bijection presented in Sec. II.2. In Sec. III.3.1, we show how the GNN performance depends on the number of snapshots used to estimate correlation functions. In Sec. III.3.2, we compare the performance of the GNN and the MLP on the snapshot data to determine the advantage of the graph preprocessing. Finally, in Sec. III.3.3, we study whether the GNN error on snapshot data scales according to the central limit theorem.
Figure 9: Performance of training cases #$\#3$ and #$\#5$ using snapshot data. A view of the metrics so far considered as a function of the cluster size. The two best cases, namely cases #$\#3$ and #$\#5$, have been retained. The dataset has been generated using a finite number of snapshots to estimate the observables, namely $N_{\text{smpl}}=1000,2500,5000,10000,25000$, and $50000$ snapshots. The error bars show the standard error on five distinct and independent trainings. The vertical gray lines indicate the point below which cluster sizes are considered in the training dataset.
III.3.1 Comparison between best training scenarios
The two best cases computed using Z-observable measurements, presented in Sec. III, are compared against each other in Fig. 9, namely the cases #$\#3$ and #$\#5$ (see Table 1). Furthermore, since when dealing with snapshots to estimate correlation functions like in experiments, one should be able to carry out measurements in the X-basis as well (see Eq. (15)), we check the GNN performance when combining Z- and X-correlators in Fig. 10. The colors were carefully chosen in Figs. 9 and 10 to match the snapshot sample size.
Looking first at Fig. 9, overall, one notices that the case #$\#5$ slightly underperforms case #$\#3$, especially for smaller $N_{\text{smpl}}$’s. The architectural benefits seem to be washed out when training with snapshot data, as is clearly apparent when looking back at the case #$\#5$ in Fig. 4. Evidently, increasing the number of snapshots is very beneficial to the training. When comparing Fig. 4 to Fig. 9, having in mind the bijection proof detailed in Sec. II.2, it seems obvious the almost-perfect rendering of the GNN (see cases #$\#3$, #$\#5$ and #$\#6$ in Fig. 4) is unlocked due to the exact and precise correlators provided, and that the main limitation seen in Figs. 9 and 10 relates to the statistically diluted physical content in snapshots.
Figure 10: Performance of the training case #$\#6$ using snapshot data. A view of the metrics so far considered as a function of the cluster size. The case #$\#6$ is considered. The dataset has been generated using the same finite number of snapshots as in Fig. 9. The error bars show the standard error on five distinct and independent trainings. Again, the gray vertical line indicates the cluster size threshold above which neural network models extrapolate. The dotted curves with corresponding colors are predictions from an MLP (including error bars).
III.3.2 Comparison between the GNN and the MLP performance
In Fig. 10, we compare the predictive errors and scaling of the GNN (full lines) and a simple MLP of the same architecture as the GNN’s task network (dashed lines). The same MLP is trained to predict an NN distance between a pair of atoms from the NN spin-spin correlation for all the atoms, disorder realizations, and cluster sizes included in the training set. The resulting MLP has a surprisingly low predictive error, apparently supported by the bijection proved in Sec. II.2. Even more surprisingly, the MLP performance is almost constant across cluster sizes and is comparable to the performance of the GNN at the largest cluster sizes.
Figure 11: Effect of the snapshot sample size $N_{\text{smpl}}$. Standard deviations (STDs) of the difference between NN relative distance predictions and targets for case #$\#6$ using snapshots (solid curves) for all cluster sizes considered. The corresponding dotted curves whose color matches that of the solid curve show how the STD would evolve with the sample size according to the central limit theorem, taking the snapshot data with $N_{\text{smpl}}=\{5000,10000,25000,50000\}$ as reference. The error bars show the standard error on five distinct and independent trainings of the PNA GNN.
III.3.3 Snapshot size effect
IV Conclusions and discussion
In this work, we have made use of the Principal Neighbor Aggregation (PNA) algorithm to predict the positions of Rydberg atomic arrays in 2D cold atom arrays. The training dataset was produced by simulating the ground state of a transverse-field Ising model using the Density Matrix Renormalization Group (DMRG) algorithm. Overall, the graph neural network (GNN) employed demonstrated a clear capacity to extrapolate outside its training domain – also the multi-layer perceptron (MLP) considered at the end of the work –, be it in terms of system sizes or shape. The GNN and MLP scalability is probably enabled from the quantum system’s side by the effective locality of the studied Hamiltonian [92]. The training input data was the local magnetization, the nearest-neighbor (NN) and next-nearest-neighbor (NNN) correlation functions, and the targets were NN and NNN relative distances between atoms in the array, depending on the training case scenarios (see Table 1). To narrow down the best training conditions for the GNN, various training procedures were followed, each with its own efficiency. The correlators were all computed in the Z measurement basis and in some cases Z- and X-correlators were used.
[…]
Based on this systematic study of the PNA GNN, we can easily identify various avenues to fine-tune the neural network and tailor it to specific needs. The points brought up in what follows should further enhance the GNN performance in predicting the NN relative distances. As a first point, i) one could add more cluster sizes to the training dataset $\{4\times 4,5\times 5,6\times 6\}$ considered here, such as sizes $4\times 5$, $5\times 6$, etc. Secondly, ii) one could increase the number of correlators to include longer-range ones, thereby amplifying the over-completion of the task of inferring NN relative distances. Thirdly, iii) one could enable higher-order graph edges that are not associated with any correlator in the datasets since this has proven to be beneficial (cases #$\#4$ and #$\#5$) to the performance; these ‘weightless’ edges would act like ‘skip connections’ in the back-propagation of the gradients during the GNN optimization. Lastly, iv) the $\vec{\Omega}$-history used to produce the training dataset could be more carefully chosen to remain near a quantum phase transition with an increased number of $\Omega$ slices.
We are currently exploring applications of the method described in this article to time-dependent experimental data on Rydberg atomic arrays. Looking forward, the GNN (and MLP) could be trained on various types of Hamiltonians [24] and, therefore, predict many possible Hamiltonian terms and parameters, e.g., identifying unknown sources of noise from the measurement data instead of only predicting parameters of a pre-defined Hamiltonian. It is also interesting to consider the limitations of the GNN scalability to larger system sizes. Eventually, one could envision feedback control of the experimental setup to reduce uncertainties on the simulated Hamiltonian.
Acknowledgements.
We are grateful to Antoine Browaeys, Thierry Lahaye, and their group at IOGS for getting us interested in this problem and for useful discussions. We thank Krzysztof Maziarz for suggesting Principal Neighborhood Aggregation as the GNN architecture. We also thank Lucy Reading-Ikkanda for the graphical design of Figs. 1 and 2. O.S. thanks Sékou-Oumar Kaba for useful discussions regarding the GNN implementations. A.D. thanks Liubov Markovich for the discussions on Hamiltonian learning techniques. O.S. acknowledges support from the Swiss National Science Foundation through the Postdoc.Mobility fellowship. This work was granted access to the HPC resources of TGCC and IDRIS under the allocation A0170510609 attributed by GENCI (Grand Equipement National de Calcul Intensif). It has also used high performance computing resources of IDCS (Infrastructure, Données, Calcul Scientifique) under the allocation CPHT 2024. This work is part of HQI initiative (www.hqi.fr) and is supported by France 2030 under the French National Research Agency award number ANR-22-PNCQ-0002. The Flatiron Institute is a division of the Simons Foundation.
Data availability
The data presented in this work is partly available on https://github.com/oliviersimard/Hamiltonian_Learning_GNN.git. The DMRG code used to produce the samples can be provided upon request to the authors. Fig. 4 can be reproduced following the instructions available through this link.
Code availability
The DMRG code used to generate the datasets is available upon request to the authors and was written using the library ITensor [94]. The PNA GNN codes to train and test upon the datasets are available on https://github.com/oliviersimard/Hamiltonian_Learning_GNN.git, as well as the datasets. The GNN code was written using utilities from PyTorch Geometric [95].
Appendix A Details on the numerical implementation and hyperparameters used
@@
For all cases laid out in Table 1, the same GNN hyperparameters were used. The number of layers defining the PNA was 4, and the number of input features that both nodes and edges corresponded to the length of the $\Omega$-history. The node embedding dimension used was 32. The edge feature dimension corresponds to the combined number of NN edges and, if applicable, NNN edges embedded into a higher dimensional space through a three-layer perceptron. 750 epochs were sufficient to train the GNN and obtain a validation loss comparable to the training one. A histogram of the degree of connectedness of the training dataset graphs is provided to the GNN during the training phase. The number of graphs in the training dataset is $2000$, times the $\Omega$-history, making it $20,000$ samples. The total number of samples is the same whether the snapshots were used to calculate the correlation function or direct evaluation was employed. The testing dataset comprises $200$ graphs times the $\Omega$-history.
@@
Appendix B Technical details of the DMRG calculations

### v2

@@ Appendix B Number of physical system sizes in training datasets
In this section of the Appendix, the effect of the number of physical system sizes included in the training datasets is studied. Here, we utilize exact training datasets.
@@
In Fig. 10, for the set of sizes $\{4\times 4,5\times 5\}$ and $\{4\times 4,5\times 5,6\times 6\}$, we plot the coefficient of determination $R^{2}$ (5) in the top panel, the MAE (6) in the middle panel, and the median of the MAE (MEDAE) in the bottom panel. The case #3\#3 is utilized. Without much surprise, the GNN performs much better at extrapolating relative NN distances $\Delta R_{\langle i,j\rangle}$ for clusters whose sizes lie outside the training set because of system size effects. It may, however, come as a surprise that the smaller system sizes, i.e. $4\times 4$ and $5\times 5$, perform worse than that $6\times 6$ within the training size domain, but it seems that this effect mostly comes from the inclusion of the $\chi^{\text{NNN}}$, as seen in Fig. 4. Since smaller clusters are subparts of larger clusters, the information of the smaller clusters aggregates and makes it possible to predict with high precision the $\Delta R$’s of clusters with sizes beyond the training set.
Figure 10: Training dataset cluster size effect. Incidence of the cluster sizes included in the exact training dataset on various metrics. In violet, clusters of sizes $4\times 4$ and $5\times 5$ were considered, whereas in light pink, one extra cluster size $6\times 6$ was considered. The vertical lines refer to the point beyond which sizes lie outside of the training range, i.e. these sizes did not take part of the training datasets. The colors of the vertical lines match with that of the curves.
Appendix C Predicting NN and NNN distances
In this section of the Appendix, we check the performance of the GNN in predicting the NN and NNN relative distances simultaneously. Again, we utilize exact training datasets.
@@ Appendix C Predicting NN and NNN distances
Regarding the cost function in Eq. (2), one could ask whether one should use both the relative NN and NNN distances whenever both $\chi^{\text{NN}}$ and $\chi^{\text{NNN}}$ are provided, or if one should use an overcomplete set of physical short-range correlators to predict only the relative NN distances, i.e. $\chi^{\text{NN}}$ and $\chi^{\text{NNN}}$. As proven in Appendix A, with arbitrary precision on the correlation functions, one would theoretically only need $\chi^{\text{NN}}$ to predict relative NN distances $\Delta R_{\langle i,j\rangle}$ between Rydberg atoms. However, we may think that giving some extra information to the GNN could be beneficial to overcome numerical or experimental imprecision. To that effect, in Fig. 11, we compare two GNN training cases #3\#3 where, on one hand, only relative NN distances are used as targets, and on the other hand, both NN and NNN relative distances are used as targets, for training sizes $\{4\times 4,5\times 5,6\times 6\}$ and $\{4\times 4,5\times 5,6\times 6,7\times 7\}$. It is very clear that the overcompletion of information helps the GNN palliate the numerical/experimental inaccuracies in determining the relative distances, meaning that predicting solely NN relative distances with $\chi^{\text{NN}}$ and $\chi^{\text{NNN}}$ at hand is better than predicting both NN and NNN relative distances with the same information throughput. It also seems that predictions of larger relative distances (NNN and above) require larger cluster sizes in the training set because edge effects prevail longer. Therefore, if one seeks to also predict NNN relative distances between atoms, one should consider training the GNN with larger cluster sizes to improve substantially the predictive power of the model (compare the navy blue curve with the pink one in Fig. 11). Interestingly, adding another task network specializing in predicting NNN distances worsened the GNN performance compared to using the same task network for predicting NN and NNN distances.
Figure 11: Target optimization schemes. Incidence on various metrics of including the NNN relative distance as a target in the cost function minimization. In blue and light pink, both the NN and NNN relative distances are predicted, while in red and pink, only the NN ones are predicted. Training sizes used are indicated in parentheses. The vertical grey lines refer to the point beyond which sizes lie outside of the training range.
Appendix D Details on the numerical implementation and hyperparameters used
@@
For all cases laid out in Table 1, the same GNN hyperparameters were used. The number of layers defining the PNA was 4, and the number of input features that both nodes and edges corresponded to the length of the $\Omega$-history. The node embedding dimension used was 32. The edge feature dimension corresponds to the combined number of NN edges and, if applicable, NNN edges embedded into a higher dimensional space through a three-layer perceptron. 550 epochs were sufficient to train the GNN and obtain a validation loss comparable to the training one. A histogram of the degree of connectedness of the training dataset graphs is provided to the GNN during the training phase. The number of graphs in the training dataset is $2000$ per physical system size, times the $\Omega$-history, making it $20,000$ samples per physical system size. The total number of samples is the same whether the snapshots were used to calculate the correlation function or direct evaluation was employed. The testing dataset comprises $200$ graphs per physical system size times the $\Omega$-history.
@@
Appendix E Technical details of the DMRG calculations